Q4) You are a fundamental analysis and you have the following data on a prospective investment

Shares in issue (millions) 2
Earnings per share (pence) 50
Dividend per share (pence) 41
Share price (pence) 542
Forecast growth rate % (constant, annualised) 3
Beta co-efficient 1.5

Other information

Average market return 8%

Gilts yield 4%

Calculate;

1. P/E ratio

2. Theoretical share price using the dividend valuation model
Answer #1

1). P/E Ratio = Share Price / EPS = 542 / 50 = 10.84 times

2). According to the CAPM,

Required Return = Risk-free Rate + [Beta * (Market Return - Risk-free Rate)]

= 4% + [1.5 * (8% - 4%)] = 4% + 6% = 10%

P0 = [D0 * (1 + g)] / [r - g]

= [41 * (1 + 0.03)] / [0.10 - 0.03]

= 42.23 / 0.07 = 603.29 pence